OpenStudy (anonymous):

4x + y = 8 2x - y = 4 a. (0,2) c. (2,0) b. (-2,0) d. (0,-2)

OpenStudy (muzzack):

lets use elimination

OpenStudy (muzzack):

4x + y = 8 2x - y = 4 ----- subtract ------ 2x=4 ----- x=2

OpenStudy (muzzack):

now subtitute 2x - y = 4 -------- 2(2) - y = 4 ----- 4-y=4 -y=0 y=0

OpenStudy (muzzack):

(2,0)

OpenStudy (anonymous):

thanks